# library.bib

### v1

@@ 3.3 Resampling, Tuning, Prediction Types, and Pre-Processing
Resampling is performed as nested cross-validation with five outer and three inner resampling folds for unbiased generalization error estimates [bischl2023hpo]. Stratified resampling is used to preserve the proportion of censoring in respective folds. The number of outer resampling folds is reduced to four if necessary to prevent folds with less than 30 observations, which affected only the veteran dataset from the survival package.
@@
Tuning is performed on the inner resampling folds of the nested cross-validation. We use random search with $50\rho$ iterations where $\rho$ is the dimensionality of the search space (ranging from 1 to 8). For example, one parameter of GLMN is tuned for 50 iterations, whereas six parameters of ORSF are tuned for 300 iterations. This method grants each method the opportunity to be tuned to the same relative amount, i.e. 50 iterations per tunable parameter. When the tuning space was finite and smaller than 50, we used exhaustive grid search to achieve the same tuning result but with lower computational cost; this applied to learners AFT (tuning across one of three distribution families), Flex (tuning $k\in 1\ldots 10$, and RRT (tuning $\mathtt{minbucket}\in 5\ldots 50$). The secondary stopping criterion for the tuning process was a time limit of 150 hours, which ensured that one outer resampling iteration (tuning and final model fitting), could complete within seven days, a constraint imposed by the computational environment. This restriction was violated most often for memory-intensive models on datasets with many observations (see Section 4 and Appendix F.3). The tuning process is repeated independently for each tuning measure (see Section 3.4).
@@
Prediction types: In general, there are four prediction types in survival analysis [sonabend2021b]: A linear predictor lp, continuous ranking crank (e.g. a relative risk), a distribution distr (e.g. the survival probability), and predicted survival times response. The response time is very uncommon and only directly provided by the Survival SVM at the time of writing. Here we focus on evaluating distribution and continuous rank predictions. However, the prediction types provided by individual methods (and implementation) can vary, which is why mlr3proba [pkgmlr3proba] compositors are used to derive missing prediction types needed for evaluation, if necessary. Where models only predict a probabilistic prediction, crank is calculated as the expected mortality derived from the distr prediction [ishwaran2008, sonabend2022chack]. When models predict only crank or lp (i.e., RRT and SSVM) then the prediction distribution is composed with a Kaplan-Meier baseline and a PH or AFT model form, which are chosen as these are common model assumptions (which, however, doesn’t seem to yield good predictions based on our results; cf. Appendix F). The XGBoost model with the Cox optimization criterion predicts lp, the distr prediction is composed using the Breslow estimator [lin2007] similarly to previous benchmarks [jaeger2024].
@@
Pre-processing is applied only if either technically required to run a model or in line with standard recommendations for that model class. This includes standardization of covariates to unit variance and zero mean and/or treatment encoding of categorical features. We created pipelines for all learners (where required) with mlr3pipelines [pkgmlr3pipelines], to combine the respective pre-processing operation with the learning algorithm, and to properly embed the pre-processing into CV. Appendix D lists the model-specific pre-processing performed. In addition to these model-specific pre-processing steps, we collapse levels of categorical variables with frequencies below 5% as part of the model pipeline, ensuring that high-cardinality categorical features are handled consistently. As we only applied basic pre-processing, no additional hyperparameters were added to the tuning search space.
@@ 3.4 Performance Evaluation
We assess performance using two primary measures alongside seven additional measures. For cases where individual model predictions were not possible during the inner- or outer resampling procedure due to any kind of error (e.g. numerical issues), the prediction of a KM estimator was used as a fallback. This ensures a statistically sound evaluation, and is a good compromise between either assigning failed models a performance of 0 (which overpenalizes) or simply disregarding failed folds during evaluation (which is overly optimistic) [mlr3booklargescale].
Measures chosen for this benchmark are summarized in Table 1. Of these measures, only three are used to provide primary results: Harrell’s C [harrell1982] for discrimination as well as the Right-Censored Log Loss (RCLL) [avati2020countdownregressiona, rindt2022survivalregression] and Integrated Survival Brier Score (ISBS) [graf1999] for overall predictive ability. The benchmark procedure is run twice, each tuning either for discrimination (Harrell’s C) or overall predictive ability (RCLL) and evaluated according to similar measures, while we use ISBS in addition to evaluate models tuned on either measure for comparison purposes. We also include the strictly proper scoring rules Re-Weighted Integrated Survival Log-Likelihood (RISLL) and Re-Weighted Negative Log Loss (RNLL) [sonabend2024scoringrules]. Furthermore, we explore the calibration measures D-Calibration [haider2020effectivewaysa] and Houwelingen’s [vanhouwelingen2000] $\alpha$ in Section G.3.
Statistical Analysis is conducted following Demšar [demsar2006], initially performing global Friedman rank sum tests for all measures, where the “groups” are the models and the “blocks” are the independent datasets. Significance after Bonferroni-Holm adjustment determines whether post-hoc tests are conducted. Post-hoc (multiple-testing corrected) Bonferroni-Dunn tests are conducted and presented as critical difference diagrams, using CPH as the reference model for comparison.
Table 1: Considered performance measures. Column 1: The measure used for evaluation. Harrell’s C, RCLL, and ISBS are used for primary analysis with remaining results in Appendix F. Column 2: The corresponding measure used for tuning, i.e., models are evaluated with (1) if they were tuned on (2). Column 3: The type of evaluation measure. Column 4: Evaluated prediction type in mlr3proba.
Evaluation Measure Tuning Measure Type Evaluates
@@
Integrated Survival Brier Score (ISBS) Harrell’s C, RCLL Scoring Rule distr
Right-Censored Log-Likelihood (RCLL) RCLL Scoring Rule distr
Re-Weighted Integrated Survival Log-Likelihood (RISLL) RCLL Scoring Rule distr
Re-Weighted Negative Log Loss (RNLL) RCLL Scoring Rule distr
D-Calibration RCLL Calibration distr
van Houwelingen’s $\alpha$ RCLL Calibration distr
@@ 3.5 Datasets
To obtain a collection of suitable datasets, we ran a search across the CRAN Task View “Survival Analysis” [^2] https://cran.r-project.org/view=Survival, Python’s pycox library and related literature (see Section 2) and existing collection of survival datasets [drysdale2022survsetopensource], yielding over 120 datasets. After applying the dataset inclusion criteria (see Section 3.1) and removing duplicates and derivations of other datasets, a total of 32 datasets remained. Minor changes are made to variable names, recoding of factor levels, and deletion of non-informative or “illegal” covariates (e.g. ID numbers). Observations are deleted if their event time is equal to zero. Since this paper is not concerned with a model’s ability to handle or impute missing covariate data, missing observations are removed, which occurred rarely and in small proportions. While this could introduce bias if missingness is dependent on the features or outcome, the goal of this study is model comparison, so this should not affect the conclusions. Lastly, in cases where datasets had a large number of unique time points, the time variable was coarsened via appropriate rounding, which greatly reduced computational coast for some methods including RAN, AK, MBST and CIF. For full details, see the pre-processing code contained in the GitHub repository (see Section 3.1). Summaries of the datasets in terms of the number of observations and covariates after modification and censoring proportions, along with citations for the respective sources, can be found in Appendix B.
@@ 4 Results
[…]
@@
Discrimination CD plots for discrimination, tuned and evaluated on Harrell’s C (Figure 1(a)), indicate that all models outperform the baselines learners (KM, NA, AK) except for RRT and GLMN. AFT and ORSF are the top-performing models but fail to significantly outperform the CPH baseline. The nine remaining models mostly belong to the classes of RSFs and GBMs in addition to Pen, which all achieve average ranks between 6.5 and 8.5, indicating similar discrimination performance.
Overall Performance CD plots for overall performance evaluated on RCLL (Figure 1(c)) indicate that only the tuned AFT significantly outperforms the CPH model. Flex ranks better than CPH but not significantly. CPH outperforms CoxB, MBST, Pen, XGBCox, and GLMN, which do not score significantly different ranks. CIF, ORSF, RAN and RSFRC rank worse here compared to the discrimination results and are significantly outperformed by CPH. Notably, NA and KM rank better on overall performance than RFSRC, AK, RRT and XGBAFT, while RAN ranks between NA and KM. Models tuned on RCLL and evaluated on ISBS (Figure 1(d)) also show CPH to be among the top performing models, in this case not significantly outperformed by any other model. CoxB achieves the best average rank, with ORSF closely behind on par with CPH. AFT here performs slightly worse yet still in the top half of all models. Pen, Flex and CIF also do not significantly differ from CPH, whereas the majority of models are outperformed by CPH, from GLMN and MBST to XGBCox, RAN, KM, RFSRC and NA. XGBAFT, AK and RRT are unambiguously placed at the bottom ranks. For models tuned with Harrell’s C and evaluated on ISBS (Figure 1(b)), we observe a similar ranking with CoxB achieving the best ranking score, followed by CPH, ORSF, Pen, Flex and CIF. RSFRC, AFT, and RAN complete the range of models not performing significantly different than CPH, whereas GLMN and the remaining GBMs, RSFs and baseline models perform significantly worse than CPH.
(a) Models tuned and evaluated with Harrell’s C
(b) Models tuned with Harrell’s C, evaluated with ISBS
(c) Models tuned and evaluated with RCLL
(d) Models tuned on RCLL, evaluated with ISBS
@@
Figure 1: Critical difference plot comparing models with the CPH reference tuned on Harrell’s C (a,b) and RCLL (c,d) and evaluated on Harrell’s C (a), RCLL (c) and ISBS (b,d). Superior models (lower ranking scores) are on the left with decreasing performance (higher rank) moving right. Models connected by thick horizontal lines are not significantly different from the baseline when adjusting for multiple comparisons.
Figure 2: Boxplots of aggregated scores across all datasets for models tuned and evaluated with RCLL showing unmodified RCLL scores (a), Explained Residual Variation (ERV) scores (b), and scores scaled such that 0 is equivalent to KM and 1 is achieved by the best model for each dataset and measure.
@@
We additionally present boxplots both for individual scores per dataset and aggregated scores. We offer three versions of these aggregated scores to support evaluation and analysis, illustrated by Figure 2: a) Raw scores as calculated by the corresponding measure; b) “Explained Residual Variation” (ERV) [korn1991] scores (similar to the “Index of Prediction Accuracy” [kattan2018indexprediction]) where negative values imply worse than KM performance, 0 is equivalent to KM, higher is better, and 1 denotes a perfect model; c) Scaled scores, whose interpretation is the same as the ERV ones, with the difference that 1 is achieved by the best model for a given task and measure [caruana2006empiricalcomparison]. Further results for all measures can be found in Appendix F. Additional tables and visualizations are available on our results website which is linked in our GitHub repository. The repository also provides downloads for all raw results and tuning archives.
@@ 5 Discussion
Discrimination The critical difference plots indicate that AFT, ORSF, and CPH were among the best performing methods, notably including two classical methods. The best-performing ML methods were ORSF ahead of CPH and RFSRC, CIF, and CoxB behind it, highlighting RSFs and GBMs good performance while not being able to significantly outperform the CPH reference. XGBoost with either Cox or AFT objective appears at the lower end of the models ranking on par with CPH, while ranking lower than the likelihood-boosting method CoxB and all RSFs. Conversely, CoxB did not require any explicit tuning outside of its internal optimization method, making it both computationally more efficient in this regard while achieving comparable or better discrimination performance. Penalized Cox (GLMN) and decision trees (RRT) are the only non-baseline models clearly outperformed by CPH, which in either case may be expected given that RRT is known to be inferior to boosted or bagged trees for prediction purposes, and GLMN representing a penalized version of CPH that may not be well suited for the low-dimensional tasks included in this benchmark. AK performs as expected, as it represents a slight improvement over KM and NA as AK takes the covariate-dependent censoring distribution into account. KM and NA are only included as a reference due to their Harrell’s C scores being constant at 0.5 by design.
Overall Performance Overall performance is judged primarily by the strictly proper scoring rule RCLL and additionally by the ISBS scoring rule [sonabend2024scoringrules] and should capture both a model’s discrimination and calibration properties. Notably both in RCLL and ISBS, the classical statistical methods (AFT, Flex, CPH) perform very well, with CoxB being the only ML approach consistently scoring among the top ranks for most evaluation metrics. CPH was only significantly surpassed by AFT when evaluated on RCLL, where AFT was superior by a wide margin. We note that our use of the parametric AFT model can be considered unconventional, as we tune the functional form (i.e. whether to use a Weibull, log-normal, or log-logistic distribution) within the three inner resampling folds, which lead to tuning results averaged across the five outer resampling folds where different functional forms were chosen by the tuning procedure. This also affects XGBAFT analogously. In real-world scenarios, a specific functional form is typically chosen in advance, leading us to believe that the performance of the parametric models may not be fully representative. We also note that many ML methods that perform on par with CPH on discrimination measures noticeably rank lower when considering overall performance measures. This implies that calibration is an important factor that may be neglected by some models, but where AFT is particularly strong. The difference in rankings between RCLL and ISBS stands out, as ORSF ranks very well on RCLL yet is significantly outranked by CPH on ISBS. Interestingly, the difference in rankings based on ISBS is very similar for models tuned on either Harrell’s C or RCLL, with CoxB, CPH and ORSF leading the ranking. While RCLL and ISBS are both scoring rules, they evaluate different properties of the predicted targets and therefore produce different results. Generally speaking, multiple measures should be considered for performance evaluation in any case as they may highlight aspects of performance relevant in different contexts.
@@
Calibration In the case of D-Calibration, a model is considered well-calibrated if the underlying Pearson’s $\chi^{2}$-test results in $p>0.05$. For van Houwelingen’s $\alpha$, calibration is indicated by values close to 1. All models aside from notable outliers AK, MBST, XGBCox and XGBAFT appear to be reasonably well calibrated based on both D-Calibration and van Houwelingen’s $\alpha$ across most datasets (Section G.3). This is consistent with the comparatively poor performance displayed in the results evaluating on RCLL and ISBS, where RRT, AK and XGBAFT were among the lowest-scoring models. However, both measures should be considered experimental, and more research on calibration-specific measures is required for a more conclusive evaluation (e.g. [austin2020graphicalcalibration]) For discrimination and overall performance, results on individual datasets are not discussed here but presented in Appendix G as the overall trend is similar to the aggregated results.

### v2

@@ 3.3 Resampling, Tuning, Prediction Types, and Pre-Processing
Resampling is performed as nested repeated cross-validation with three outer and three inner resampling folds for unbiased generalization error estimates (Bischl et al., 2023). Outer cross-validation is repeated five times for datasets with a total number of events in $[500,2500)$ and ten times for datasets with fewer than 500 events. This aides the stability of performance estimates by increasing the total number of evaluations to up to 30 (10 repetitions $\times$ 3 outer folds) for the datasets with few events, while saving computational costs for the largest tasks where the available training data is likely sufficient for reliable evaluation. Inner cross-validation is always repeated twice, leading to six total evaluations per tuning iteration. Resampling is stratified by event indicator to preserve the censoring proportion in respective folds.
@@
Tuning is performed on the inner resampling folds of the nested cross-validation. We use Bayesian optimization (Garnett, 2023; Bischl et al., 2023) with $50\rho$ iterations, where $\rho$ is the dimensionality of the search space, ranging from 1 to 8. For example, one parameter of GLMN is tuned for 50 iterations, whereas six parameters of ORSF are tuned for 300 iterations. This method grants each method the opportunity to be tuned to the same relative amount: 50 iterations per tunable parameter. When the tuning space was finite and smaller than 50, we exhausted all configurations to achieve the same tuning result but with lower computational cost; this applied to learners AFT (tuning across three distribution families), Flex (tuning $k\in\{2,\ldots,10\}$, and RRT (tuning $\mathtt{minbucket}\in\{5,6,\ldots,50\}$). We set a time limit for the tuning process of 150 hours to ensure that one outer resampling iteration (tuning and final model fitting) could be completed within seven days, a constraint imposed by the computational environment. This restriction was most frequently violated for memory-intensive models on datasets with many observations, where in some cases the final evaluation was unsuccessful (see Section˜4 and Appendix E.3). The tuning process is repeated independently for each tuning measure (see Section˜3.4).
@@
Prediction types: In general, there are four prediction types in survival analysis (Sonabend, 2021): A linear predictor lp, continuous ranking crank (e.g., a relative risk), a distribution distr (e.g., the survival probability), and predicted survival times response. The response time is very uncommon due to its generally poor quality (Henderson et al., 2001) and only directly provided by the Survival SVM and XGBAFT at the time of writing. We focus on evaluating distribution and continuous rank predictions. The prediction types provided by individual methods (and implementations) can vary, which is why mlr3proba (Sonabend et al., 2021) compositors are used to derive missing prediction types needed for evaluation when appropriate. Where models only predict a probabilistic prediction, crank is calculated as the expected mortality derived from the distr prediction (Ishwaran et al., 2008; Sonabend et al., 2022). When models predict only crank or lp, then we exclusively evaluate them on discrimination measures, which is the case for RRT, XGBAFT, MBSTAFT, and SSVM. The XGBCox predicts lp, and the distr prediction is composed using the Breslow estimator (Lin, 2007) similarly to previous benchmarks (Jaeger et al., 2024).
@@
Pre-processing is applied only if either technically required to run a model or in line with standard recommendations for that model class. This includes standardization of covariates to unit variance and zero mean and/or dummy encoding of categorical features. We created pipelines for all learners (where required) with mlr3pipelines (Binder et al., 2021), to combine the respective pre-processing operation with the learning algorithm, and to properly embed the pre-processing into the cross-validation. Appendix C lists the model-specific pre-processing performed. In addition to these model-specific pre-processing steps, we collapse levels of categorical variables with frequencies below 5% as part of the model pipeline, ensuring that high-cardinality categorical features are handled consistently. As we only applied basic pre-processing, no additional hyperparameters were added to the tuning search space.
@@ 3.4 Performance Evaluation
We assess performance using two primary measures alongside four additional measures. For cases where individual model predictions were not possible during the inner- or outer resampling procedure due to any kind of error (computational issues, non-convergence, etc.), the prediction of the KM estimator was used as a fallback result. This ensures a statistically sound evaluation, and is considered a reasonable compromise between either overpenalizing models by inserting some constant value which would depend on the given metric and dataset, or simply disregarding failed iterations during evaluation, which is overly optimistic (see e.g. Fischer et al., 2024).
Measures chosen for this benchmark are summarized in Table˜1. Of these measures, only two are used to provide primary results: Harrell’s C (Harrell et al., 1982) for pure discrimination and the Integrated Survival Brier Score (ISBS) (Graf et al., 1999) for overall predictive ability, including calibration. The benchmark procedure is run twice, tuning either for discrimination (Harrell’s C) or overall predictive ability (ISBS) and evaluated according to similar measures. For integrated measures, we use the commonly chosen 80% quantile as the upper bound (Sonabend et al., 2025). We additionally explore the calibration measures D-Calibration (Haider et al., 2020) and van Houwelingen’s $\alpha$ (van Houwelingen, 2000), which we only apply on models tuned with ISBS.
Statistical Analysis is conducted following Demšar (Demšar, 2006), initially performing global Friedman rank sum tests for all measures, where the “groups” are the models and the “blocks” are the independent datasets. Significance after Bonferroni-Holm adjustment determines whether post-hoc tests are conducted. Post-hoc (multiple-testing corrected) Bonferroni-Dunn tests are conducted and presented as critical difference diagrams, using CPH as the reference model for comparison.
Table 1: Performance measures used in the benchmark, including their type and the quantity they evaluate, i.e., continuous ranks (crank) or a survival distribution (distr). Harrell’s C and ISBS are used for primary analysis with remaining results in Appendix E. Models evaluated with each measure were tuned on the corresponding tuning measure.
Measure Tuned on Type Evaluates
@@
ISBSa ISBS Scoring Rule distr
ISLLb ISBS Scoring Rule distr
D-Calibration ISBS Calibration distr
$\alpha$-Calibrationc ISBS Calibration distr
aIntegrated Survival Brier Score bIntegrated Survival Log-Likelihood cvan Houwelingen’s $\alpha$
@@ 3.5 Datasets
To obtain a collection of suitable datasets, we ran a search across the CRAN Task View “Survival Analysis” [^2] https://cran.r-project.org/view=Survival, Python’s pycox library and related literature (see Section˜2) and existing collection of survival datasets (Drysdale, 2022), yielding over 120 datasets. After applying the dataset inclusion criteria (see Section˜3.1) and removing duplicates and derivations of other datasets, a total of 34 datasets remained. Minor changes are made to variable names, recoding of factor levels, and deletion of non-informative or “illegal” covariates like ID numbers. Observations are deleted if their event time is equal to zero. Since this benchmark is not concerned with a model’s ability to handle or impute missing covariate data, observations with missing values were removed, which occurred very rarely. Lastly, in cases where datasets had a large number of unique time points, the time variable was coarsened by appropriate rounding, greatly reducing computational cost for some methods, including RAN, AK, and CIF. For full details, see the pre-processing code contained in the GitHub repository (see Section˜3.1). Summaries of the datasets in terms of the number of observations and covariates after modification and censoring proportions, along with citations for the respective sources, can be found in Appendix B.
@@ 4 Results
Global Friedman tests were significant for all measures, indicating the presence of significant differences between models and allowing for post-hoc analysis. The number of times models failed to compute results due to either time or memory constraints and required the score imputation using KM (Section˜3.3) is tabulated in Appendix E.3. We present critical difference (CD) plots for the baseline-comparison to CPH for both discrimination and overall performance (see Section˜3.4, Demšar (2006)). The top line of a CD plot represents a model’s average performance rank across all datasets in the benchmark, where a lower ranking implies better performance regardless of the evaluation measure applied. Thick horizontal lines around the CPH model rank indicate the symmetric critical difference, meaning that other models within this range do not significantly differ in rank from the reference model.
4.1 Discrimination
@@
CD plots for discrimination, tuned and evaluated on Harrell’s C (Figure˜1, top), indicate that all models outperform the baseline learners (KM, NA, AK) as expected. MBSTAFT, AFT, RAN, and CoxB are the top-performing models but fail to significantly outperform the CPH baseline. The remaining models mostly belong to the classes of RSFs and GBMs, which all achieve average ranks between 6.5 and 8, indicating similar discrimination performance. Pen, XGBCox, XGBAFT, and Flex fall slightly behind, but still lie within the critical difference.
4.2 Overall Performance
The CD plot for overall performance tuned and evaluated on ISBS (Figure˜1, bottom) similarly indicates that no model significantly outperforms CPH. Only ORSF and CoxB perform better, while CIF ranks almost identically to CPH. Pen, MBSTCox, RAN, and RFSRC fall slightly behind, with AFT and Flex further behind. XGBCox and GLMN are the only non-baseline models significantly outperformed by CPH.
@@ 4.2 Overall Performance
Figure 1: Critical difference plots comparing models with the CPH reference tuned and evaluated on either Harrell’s C (top) or ISBS (bottom). Superior models (lower ranking scores) are on the left with decreasing performance (higher rank) moving right. Models connected by thick horizontal lines are not significantly different from the baseline when adjusting for multiple comparisons.
@@
We additionally present boxplots both for individual scores per dataset and aggregated scores. We offer three versions of these aggregated scores to support evaluation and analysis, illustrated by Figure˜2: a) Raw scores as calculated by the corresponding measure; b) “Explained Residual Variation” (ERV) (Korn and Simon, 1991) scores similar to the “Index of Prediction Accuracy” (Kattan and Gerds, 2018) where negative values imply performance worse than KM, 0 is equivalent to KM, and 1 denotes a perfect model; c) Scaled scores, whose interpretation is the same as the ERV ones, with the difference that 1 is achieved by the best model for a given task and measure (Caruana and Niculescu-Mizil, 2006).
Figure 2: Boxplots of aggregated scores across all datasets for models tuned and evaluated with ISBS showing unmodified ISBS scores (a), Explained Residual Variation (ERV) scores (b), and scores scaled such that 0 is equivalent to KM and 1 is achieved by the best model for each dataset and measure. AK is omitted due to largely off-the-scale negative ERV values to avoid scaling issues.
4.3 Calibration
Calibration results presented in this section, measured by D-Calibration and van Houwelingen’s $\alpha$ (see Table 1), were obtained by tuning all considered model for ISBS. We hope that results for these measures provide additional insights, but would like to point out that they might be considered somewhat experimental with respect to underlying theory and respective implementation.
D-Calibration considers a model well-calibrated if the p-value of the underlying test is greater than 0.05, which we display in the form of a heatmap with X indicating a significant test result, indicating a model is not well-calibrated. Here, NEL, RFSRC, ORSF, and RAN are among the best calibrated, while AK, XGBCox, CIF, AFT and CPH would be considered poorly calibrated.
Figure 3: D-Calibration p-value heatmap across all datasets and models. ‘X’ indicates $p<0.05$ while a non-significant result indicates good calibration.
Van Houwelingen’s $\alpha$ relates predicted and observed hazards, with a value close to 1 implying a well-calibrated model. Here, KM, NEL, GLMN, and Pen show good calibration, while XGBcox shows a wide spread indicating poor calibration. RFSRC, RAN, and ORSF similarly show more skewed scores, indicating poor calibration.
Figure 4: Calibration scores across all datasets for each model using van Houwelingen’s $\alpha$ indicate good calibration when values are close to 1.
Further results for all measures can be found in Appendix E. Additional tables and visualizations are available on our results website which is linked in our GitHub repository. The repository also provides downloads for all raw results, including tuning archives for the close to 20.000 individual tuning procedures performed in this experiment.
@@ 5 Discussion
Discrimination The critical difference plots show that MBSTAFT, AFT, RAN, CoxB, and CPH were the five highest-scoring methods, notably including two classical methods. The best-ranking ML methods both used the AFT model form but do not significantly outperform CPH. XGBoost with either Cox or AFT objective appears at the lower end of the models’ ranking, on par with CPH, while ranking lower than the likelihood-boosting method CoxB and all RSFs. Conversely, CoxB did not require any explicit tuning outside of its internal optimization method, making it computationally more efficient while achieving comparable or better discrimination performance. Penalized Cox (GLMN), decision trees (RRT) and the SSVM are the only non-baseline models outperformed by CPH, which in either case may be expected given that RRT is known to be inferior to boosted or bagged trees for prediction purposes, and GLMN represents a penalized version of CPH that may not be well suited for the low-dimensional tasks included in this benchmark, while Pen, on the other hand, is a similar method that did rank better by some margin. AK performs as expected, representing a slight improvement over KM and NA. KM and NA are included solely as a reference because their Harrell’s C scores are intentionally constant at 0.5. We note that our use of the parametric AFT model can be considered unconventional, as we tune the distribution family (i.e., whether to use a Weibull, log-normal, or log-logistic distribution) within the tuning iterations, thus performance evaluation in the outer resampling iterations could be based on a different distribution. This also affects XGBAFT analogously. In real-world scenarios, a specific distribution is typically chosen in advance.
Overall Performance Overall performance is judged by the ISBS scoring rule (Sonabend et al., 2025) and should capture both a model’s discrimination and calibration ability. We can only evaluate models on ISBS which provide a survival distribution prediction, therefore omitting RRT, XGBAFT, MBSTAFT, and SSVM in this setting as described in Section 3.3.
CoxB again performs quite well, indicating it is an effective method in both settings. CPH in third place shows its general effectiveness, being on par with the computationally more complex CIF. Pen again outperforms GLMN by a wide margin despite their similarity, and analogously, the GBMs with Cox-objective MBSTCox and XGBCox are also separated by more than three places in ranking. The main difference between GLMN and Pen is that the former performs internal cross-validation for its primary regularization parameter, which in this setting could present a disadvantage. For XGBCox, it should be noted that this method encountered many computational errors, leading to imputed performance values as described in Section 3.4. While the AFT-based approaches seemed to outrank the Cox-based ones on discrimination, here we see AFT falling behind CPH and other Cox-based methods. The AK baseline shows noticeably worse performance than NEL and KM baselines, while it outperformed them in the discrimination setting. Methods besides CPH that performed well using either evaluation metric include CoxB, ORSF, and MBSTCox. Since not all methods could be evaluated on both metrics, the critical difference in each comparison is affected by the differing number of methods, complicating a direct comparison between the two resulting rankings. We also note that many ML methods that perform on par with CPH on discrimination measures noticeably rank lower when considering overall performance measures like ISBS.
@@
Calibration In the case of D-Calibration, a model is considered well-calibrated if the underlying Pearson’s $\chi^{2}$-test results in $p>0.05$. For van Houwelingen’s $\alpha$, calibration is indicated by values close to 1. All models, aside from notable outlier AK, as well as XGBCox and CIF, appear to be reasonably well calibrated based on D-Calibration, while results seem to depend largely on the dataset. For van Houwelingen’s $\alpha$, XGBCox is the least calibrated model apart from AK, which we omitted from the plot due to its greatly out-of-range scores. This is consistent with the comparatively poor performance displayed in the results evaluating on ISBS, where both AK and XGBCox scored low. We note that both calibration measures should be considered experimental, and more research on calibration-specific measures will be required for a more conclusive evaluation (e.g. Austin et al. (2020)) Generally speaking, multiple measures should be considered for performance evaluation, as they may highlight aspects of performance relevant in different contexts. For discrimination and overall performance, results on individual datasets are not discussed here but presented in Appendix F as the overall trend is similar to the aggregated results, with expected variation between individual tasks.